# A fixed phase tunable directional coupler based on coupling tuning

### v1

@@ I Introduction
To meet the demands of more advanced applications, the integration density of PICs is increasing rapidly [2]. The most adaptable strategies involve constructing waveguide mesh structures or multi-port universal interferometers [11, 12, 13] based on the fundamental tunable device that realizes universal 2$\times$2 transformation. This device is implemented by the Mach-Zehnder interferometer (MZI) with an extra phase shifter for relative phase tuning of the output states [6, 7, 14, 15, 16, 17]. Such an MZI consists of two balanced (50:50) beam splitters and a phase shifter for reflectivity tuning (Fig 1 (a) and (d)). However, employing a one-to-one component mapping approach to transferring MZIs from bulk free-space optics to PICs may not be the optimal strategy. This is because the design of MZIs is sensitive to fabrication errors and exhibits more bending losses [18, 19]. Achieving balanced directional couplers (DCs) requires perfect fabrication and is crucial for constructing MZIs to achieve 0-1 reflectivity. Additionally, each MZI has 4 bending sections (Fig 1 (a)), which introduces optical losses and can decrease the fidelity of larger scale circuits [12].
@@
In this paper, we propose a $\Delta{C}$-based TDC ($\Delta{C}$-TDC) (see the illustration in Fig 1 (c) and the device schematic Fig 2 (c)) that implements the same unitary as an MZI, but possesses fewer bending sections and enhanced fabrication error tolerance, with a method to assess its performance. This addresses the scalability challenges, potentially enabling more advanced applications associated with using MZIs. Unlike conventional $\Delta\beta$-TDCS, our $\Delta{C}$-TDC design achieves tunability through coupling tuning, which maintains the relative phase between coupled waveguides fixed (Fig 1 (e) and (f)) and is significantly less sensitive to coupling length and operating wavelength. Moreover, our design eliminates the issue of non-even responses from controllers in the dual-drive design [23, 24], ensuring consistent and reliable performance over an extra-wide range of wavelengths.
@@ II.1 Geometry design
A $\Delta{C}$-TDC exhibits tunability due to the variation of the coupling coefficient in the coupled region (Fig 3) while ensuring that the odd mode is perfectly antisymmetric and the even mode perfectly symmetric, thereby resulting in equal propagation constants for both waveguides [29]. To achieve this uniformity in propagation constants, it is imperative to maintain identical refractive index profiles for the coupled waveguides.
@@ II.2 Performance optimization
where the waveguide propagation constants are assumed to be identical with their values derived from the supermodes $\beta_{l}(V,\lambda)$=$\beta_{r}(V,\lambda)$ [37]. The unitary of the $\Delta{C}$-TDC is given by the time evolution of the Hamiltonian $H(V,\lambda)$ over the effective coupled length $L$
@@
The unitary of a DC can be written in the form of [38]
@@ II.3 Proposed Devices
II.4 Fabrication error analysis
TFLN is a relatively new technology, and accordingly, the fabrication process is somewhat immature [3]. We analyze how fabrication errors affect the tunability of our couplers with respect to film thickness, gap distance, etching depth, and waveguide top width at a wavelength of 1.55 $\mu$m as shown in Fig 6 with a fixed 60 sidewall angle. The sidewall angle is given by the chemistry settings of the etching process and can be kept constant. We postulate a uniform fabrication quality across both waveguides within the coupler structure while addressing individual error sources sequentially. The simulation results indicate that a $\sim$100 nm fabrication error does not significantly affect the tunability and larger errors ($\sim$200 nm) can be compensated by increasing the coupler length or operating voltage range.
III Discussion
@@
Table 1: Comparison of fundamental reflectivity tuning devices. [^2] MZI and $\Delta\beta$-DC can only operate at specific wavelengths in a periodic manner that allow them to achieve full 0-1 reflectivity, whereas $\Delta{C}$-DCs have continuous wide operational wavelength bandwidth.
@@
Several crucial aspects need to be considered when comparing the design of the three fundamental reflectivity-tuning devices. A summarized comparison of fundamental tuning devices is presented in Table 2. In terms of fabrication error, MZIs are well-known for being sensitive because of the requirement of perfect balanced directional couplers to achieve a full 0-1 reflectivity. Additionally, directional couplers are designed for specific operating wavelengths, rendering MZIs wavelength-dependent devices. On the other hand, $\Delta{\beta}$-TDCs have to be designed to present 0 reflectivity when static to ensure a complete 0-1/1-0 reflectivity, which makes such devices wavelength-dependent and less resilient to fabrication errors. On the contrary, $\Delta{C}$-TDCs can achieve a complete 0-1/1-0 reflectivity in a wide range of wavelengths and have shown robustness to fabrication error in simulations. Regarding the independent phase control required in applications such as quantum information processing, all three devices need an extra phase shifter. MZIs and $\Delta{C}$-TDCs can maintain a constant phase when working as reflectivity tuning devices, resulting in straightforward phase control. However, the phase of a $\Delta{\beta}$-TDC changes while its reflectivity is tuned and requires device-dependent characterization, which is challenging in practice. As for the driving voltage, $\Delta{C}$-TDCs require higher amplitudes compared to MZIs and $\Delta{\beta}$-TDCs when considering the same device size given the state-of-the-art technologies [20, 43]. This is because propagation constants tuning is more efficient for MZIs and $\Delta{\beta}$-TDCs. However, the driving voltage for $\Delta{C}$-TDCs can be decreased by increasing the length of the coupled region, decreasing the gap between electrodes, or better materials. Finally, in terms of the number of required bending sections, an MZI needs two more sections than both types of TDCs. This causes more losses, which becomes problematic when cascading multiple stages of MZIs [18, 49, 50]. Such cascading designs aim to overcome the challenges associated with the need for ideal balanced DCs in the MZIs. TDCs require fewer bendings, which makes them potentially experience less photon loss in large-scale circuits [11, 12, 13]. However, they may need to be designed longer to compensate for fabrication errors and high driving voltage requirements.
@@
In addition to Z-cut TFLN, our design can be implemented with X-cut TFLN, although more challenging. This is due to the requirements of the electrode positioning [51] and the need for a small waveguide geometry that can only be realized with cutting-edge thin film technology. While nanofabrication technology in LNOI has facilitated the development of numerous wafer-scale devices, it may not be sufficient to meet the requirements for such devices. Enhancing sidewall roughness and designing efficient mode size converters are additionally required. These will contribute to minimizing the optical losses due to the small patterning features, as well as ensuring the practical viability of the proposed device.

### v3

@@ Introduction
To meet the demands of more advanced applications, the integration density of PICs is increasing rapidly [2]. The most adaptable strategies involve constructing waveguide mesh structures or multi-port universal interferometers [11, 12, 13] based on the fundamental tunable device that realizes universal 2$\times$2 transformation. This device is implemented by the Mach-Zehnder interferometer (MZI) with an extra phase shifter for relative phase tuning of the output states [6, 7, 14, 15, 16, 17]. Such an MZI consists of two balanced (50:50) beam splitters and a phase shifter for reflectivity tuning (Fig 1 (a) and (d)). (Reflectivity $\eta$ is defined as $\eta=1-\kappa$, following the convention in the field of quantum optics [18, 19], where $\kappa$ represents the coupling or power transfer ratio [20, 21].) However, employing a one-to-one component mapping approach to transferring MZIs from bulk free-space optics to PICs may not be the optimal strategy. This is because the design of MZIs is sensitive to fabrication errors and exhibits more bending losses [22, 23]. Achieving balanced directional couplers (DCs) requires perfect fabrication and is crucial for constructing MZIs to achieve 0-1 reflectivity. Additionally, each MZI has 4 bending sections (Fig 1 (a)), which introduces optical losses and can decrease the fidelity of larger scale circuits [12].
@@
In this paper, we propose a $\Delta{C}$-based TDC ($\Delta{C}$-TDC) (see the illustration in Fig 1 (c) and the device schematic Fig 2 (c)) that implements the same unitary transformation as an MZI, but possesses fewer bending sections and enhanced fabrication error tolerance, with a method to assess its performance. This addresses the scalability challenges, potentially enabling more advanced applications associated with using MZIs. Unlike conventional $\Delta\beta$-TDCS, our $\Delta{C}$-TDC design achieves tunability through coupling tuning, which maintains the relative phase between coupled waveguides fixed (Fig 1 (e) and (f)) and is significantly less sensitive to coupling length and operating wavelength. Moreover, our design eliminates the issue of non-even responses from controllers in the dual-drive design [26, 27], ensuring consistent and reliable performance over an extra-wide range of wavelengths.
Results and methods
@@
A $\Delta{C}$-TDC exhibits tunability due to the variation of the coupling coefficient in the coupled region (Fig 3 (a) and (b)) while ensuring that the odd mode is perfectly antisymmetric and the even mode perfectly symmetric, thereby resulting in equal propagation constants for both waveguides [20]. To achieve this uniformity in propagation constants, it is imperative to maintain identical refractive index profiles for the coupled waveguides.
@@
where the waveguide propagation constants are assumed to be identical with their values derived from the supermodes $\beta_{l}(V,\lambda)$=$\beta_{r}(V,\lambda)$ [39].
The unitary transformation implemented by the $\Delta{C}$-TDC is given by the time evolution of the Hamiltonian $H(V,\lambda)$ over the effective coupled length $L$
@@
The unitary transformation implemented by a DC can be written in the form of [40]
@@
Additionally, when the gap between two waveguides becomes very small, it is necessary to consider the effects of higher-order modes, including radiation modes, as they can introduce losses and limit the efficiency of power transfer. However, our waveguide structure is designed to support only the propagation of fundamental modes (single-mode waveguide), and the gap is not small enough for the two waveguides to behave as a single waveguide. Therefore, the impact of higher-order modes on the overall power transfer ratio is negligible for our designs, which is verified in 3D Finite-difference time-domain (FDTD) simulations, as shown in the supplementary materials.
@@
Fabrication error analysis
TFLN is a relatively new technology, and accordingly, the fabrication process is somewhat immature [3]. We analyze how fabrication errors affect the tunability of our couplers with respect to film thickness, gap distance, etching depth, and waveguide top width at a wavelength of 1.55 $\mu$m as shown in Fig 6 (a), (b), (c) and (d) respectively, with a fixed 60 sidewall angle. The sidewall angle is given by the chemistry settings of the etching process and can be kept constant. We postulate a uniform fabrication quality across both waveguides within the coupler structure while addressing individual error sources sequentially. The simulation results indicate that a $\sim$100 nm fabrication error does not significantly affect the tunability and larger errors ($\sim$200 nm) can be compensated by increasing the coupler length or operating voltage range.
In addition to symmetric fabrication errors discussed above, asymmetric fabrication errors may also occur in practice. Such errors can result in dimensional differences between waveguides, preventing the device from achieving 0-1/1-0 reflectivity due to phase mismatches [20]. This challenge can be mitigated by improving fabrication quality or by introducing additional tuners to locally adjust the propagation constants of individual waveguides, similar to the dual-drive design [27]. This approach can be feasible in electro-optic devices, as electro-optic effects offer precise control over local parameters [8, 51].
Discussion
@@
Table 1: Comparison of fundamental reflectivity tuning devices. (${}^{\text{a}}$ MZI and $\Delta\beta$-DC can only operate at specific wavelengths in a periodic manner that allow them to achieve full 0-1 reflectivity, whereas $\Delta{C}$-DCs have continuous wide operational wavelength bandwidth.)
@@
Several crucial aspects need to be considered when comparing the design of the three fundamental reflectivity-tuning devices. A summarized comparison of fundamental tuning devices is presented in Table 1. In terms of fabrication error, MZIs are well-known for being sensitive because of the requirement of perfect balanced directional couplers to achieve a full 0-1 reflectivity. Additionally, directional couplers are designed for specific operating wavelengths, rendering MZIs wavelength-dependent devices. On the other hand, $\Delta{\beta}$-TDCs have to be designed to present 0 reflectivity when static to ensure a complete 0-1/1-0 reflectivity, which makes such devices wavelength-dependent and less resilient to fabrication errors. On the contrary, $\Delta{C}$-TDCs can achieve a complete 0-1/1-0 reflectivity in a wide range of wavelengths and have shown robustness to fabrication error in simulations. Regarding the independent phase control required in applications such as quantum information processing, all three devices need an extra phase shifter. MZIs and $\Delta{C}$-TDCs can maintain a constant phase when working as reflectivity tuning devices, resulting in straightforward phase control. However, the phase of a $\Delta{\beta}$-TDC changes while its reflectivity is tuned and requires device-dependent characterization, which is challenging in practice. As for the driving voltage, $\Delta{C}$-TDCs require higher amplitudes compared to MZIs and $\Delta{\beta}$-TDCs when considering the same device size given the state-of-the-art technologies [21, 45]. This is because propagation constants tuning is more efficient for MZIs and $\Delta{\beta}$-TDCs. However, the driving voltage for $\Delta{C}$-TDCs can be reduced by increasing the length of the coupled region. Alternatively, the figure of merit (FOM) can be improved by decreasing the gap between electrodes or by utilizing materials with stronger electro-optic effects, potentially in a hybrid fashion [52]. Finally, in terms of the number of required bending sections, an MZI needs two more sections than both types of TDCs. This causes more losses, which becomes problematic when cascading multiple stages of MZIs [22, 53, 54]. Such cascading designs aim to overcome the challenges associated with the need for ideal balanced DCs in the MZIs. TDCs require fewer bendings, which makes them potentially experience less photon loss in large-scale circuits [11, 12, 13]. However, they may need to be designed longer to compensate for fabrication errors and high driving voltage requirements.
@@
In addition to Z-cut TFLN, our design can be implemented with X-cut TFLN, although more challenging. This is due to the requirements of the electrode positioning [55] and the need for a small waveguide geometry that can only be realized with cutting-edge thin film technology. While nanofabrication technology in LNOI has facilitated the development of numerous wafer-scale devices, it may not be sufficient to meet the requirements for such devices. Moreover, based on state-of-the-art studies [10, 22, 23, 29, 17], it is reasonable to assume that each bending section contributes at least 0.1 dB loss. Additionally, TFLN waveguides with polished surface roughness have demonstrated propagation loss as low as 0.027 dB/cm [56]. Given that the $\Delta{C}$-TDC has one fewer bending section, a device no longer than approximately 3.7 cm with polished waveguides would have a loss advantage over MZIs that contain two bending sections. Enhancing sidewall roughness as well as designing efficient mode size converters are required, as these improvements will contribute to minimizing optical losses due to the small patterning features and ensuring the practical viability of the proposed device.